# Improving Multi-candidate Speculative Decoding

### v1

@@ 4.1 Experimental Setup
4.2 Overall Results
Figure 4: Speedup ratios compared to vanilla inference for different SD methods based on MT-Bench dataset using greedy sampling. The bars represent speedup ratios for different model combinations (draft model + target model)
Figure 5: Left graph shows the relationship between width and acceptance rate ($\alpha$) and right graph shows the the relationship between width and generation speed in tokens/s
Dataset Methods Configuration LLaMA 2-7B Chat Vicuna-7B
Alpaca Baseline SD $\gamma=5$ 0.45 0.43
MCSD $4\times 2\times 2\times 1\times 1$ 0.64 0.63
Our method $D=5,W=16$ 0.92 0.94
TriviaQA Baseline SD $\gamma=5$ 0.56 0.42
MCSD $4\times 2\times 2\times 1\times 1$ 0.72 0.59
Our method $D=5,W=16$ 0.96 0.94
MT-Bench Baseline SD $\gamma=5$ 0.43 0.49
MCSD $4\times 2\times 2\times 1\times 1$ 0.63 0.68
Our method $D=5,W=16$ 0.93 0.93
Table 2: Comparison of acceptance rate ($\alpha$) for different methods using TinyLlama-1.1B as draft model with greedy sampling (temp = 0)
The experiment result shown in Figure 4 demonstrates substantial improvements in generation speed. All of the configurations are shown and explained in Section 4.1. Conducted on the MT-Bench dataset, our method achieves a maximum of 2.96 times speedup over the baseline, mainly due to the target model initialized multi-candidate token tree.
Also, we measure the effect of width on acceptance rate and speed in Figure 5. As the width increases, we consistently observe an improvement in acceptance rates and speed across different model pairs. The acceptance rate curve converges when width=12, and the speed curve converges in width=10. These findings demonstrate our method’s effectiveness in improving acceptance rates and speed with relatively small tree widths. In addition, we notice there is a noticeable drop in speed when width=21. This speed decline is due to the increased computational overhead associated with processing a more significant number of candidate tokens in parallel, which begins to outweigh the benefits of speculative decoding.
4.3 Acceptance Rate
@@
4.4 Ablation Study
Figure 6: Both graphs show the relationship between width and generation speed. Draft model on the left graph: Llama-68M. Draft model on the right graph: TinyLlama-1.1B
To show the effect of the decision model in our framework, we compare the inference speeds under two scenarios, using or not using the decision model with different parameter size draft models. We still use MT-bench as input prompts, and all other settings remain identical. As a result, in Figure 6, both large and small draft models return higher speeds without using the decision model most of the time. The decision model only improves the speed slightly when the width is low. However, in either case, the speed improvement of the decision model is minimal. In addition, it proves that our method’s speedup is mainly due to the target model’s initialized multi-candidate token tree and topology-aware causal mask rather than the decision model.
5 Conclusions
In this work, we proposed a target model initialized multi-candidate token tree that successfully increases the acceptance rate for multi-candidate speculative decoding and a dynamic mask slicing technique to avoid topology-aware causal mask generation overhead for dynamic multi-candidate speculative decoding. While we have yet to discover the decision model that could make dynamic multi-candidate speculative decoding faster than static one under all scenarios, our work can help future research to find a better decision model and process for dynamic speculative decoding.

### v2

@@ A.1 Dynamic MCSD Configurations and Results
A.1.1 Effects of Width
Figure 6: Left graph shows the relationship between width and acceptance rate ($\alpha$) and right graph shows the the relationship between width and generation speed in tokens/s under MCSD dynamic configuration
We measure the effect of width on acceptance rate and speed in Figure 6 with fixed depth $D=5$ based on empirical results. As the width increases, we consistently observe an improvement in acceptance rates and speed across different model pairs. The acceptance rate curve converges when $W=12$, and the speed curve converges in $W=10$. These findings demonstrate our method’s effectiveness in improving acceptance rates and speed with relatively small tree widths. In addition, we notice there is a noticeable drop in speed when $W=21$. This speed decline is due to the increased computational overhead associated with processing a more significant number of candidate tokens in parallel, which begins to outweigh the benefits of speculative decoding.
A.1.2 Acceptance Rate
@@
A.2 Decision Model
Figure 7: This figure demonstrates that, when using hidden states as input with labels of 0 or 1, the three-layer neural network struggles to effectively distinguish between negative and positive examples. As a result, a significant area of overlap is observed in the predictions.
Figure 8: The left figure compares the acceptance rate, whether or not the decision is used, and the right figure compares its speed.
To show the effect of the decision model in our framework, we compare the inference speeds under two scenarios, using or not using the decision model with different parameter size draft models. We still use MT-bench as input prompts, and all other settings remain identical. As a result, in Figure 8, both large and small draft models return higher speeds without using the decision model most of the time. The decision model only improves the speed slightly when the width is low. Nonetheless, in either case, the speed improvement of the decision model is minimal. In addition, it proves that our method’s speedup is mainly due to the target model’s initialized multi-candidate token tree and topology-aware causal mask rather than the decision model.